Ok so here is my word problem:
A university bookstore recently sold a wirebound graph-paper notebook for $1.64, and a college-ruled notebook for $1.17. At the start of spring semester, a combination of 50 of these notebooks were sold for a total of $66.02. How many of each type were sold?
So far what I have is:
x+y=50
1.64x+1.17y=66.02
I found that x=50-y but now I am completely lost. Please help. Thanks in advance.

Answer by Alan3354(30993)

A university bookstore recently sold a wirebound graph-paper notebook for $1.64, and a college-ruled notebook for $1.17. At the start of spring semester, a combination of 50 of these notebooks were sold for a total of $66.02. How many of each type were sold?
So far what I have is:
x+y=50
1.64x+1.17y=66.02
I found that x=50-y but now I am completely lost. Please help. Thanks in advance.
=============
That's a good start.
Sub for x in the 2nd equation
1.64x+1.17y=66.02
Eliminate fractions first
164x+117y=6602
164*(50-y) + 117y = 6602
8200 - 164y + 117y = 6602
-47y = - 1598
y = 34 = 34 college-ruled, whatever that is.
-----------
x = 50-y = 16 graph-paper thingies